|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
|
|
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
|
|
SASE1 Status and SASE3 Readiness
Technical integration and coordination
|
| Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 447
Entry time: 30 Aug 2023, 10:55
|
| Author: |
Frederik Wolff-Fabris |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Notes XO Integration Meeting 30.08.2023 |
|
|
General Announcement:
- Facility development week is ongoing.
- After the cryo incident last weekend the machine colleagues have an impression everything is in good shape and next tuning day can be made on Monday.
- Next UOM will have an invitation for instruments to briefly report on jitter issues (pointing, temporal, etc) that impact performance from last weeks.
If you have found any jitter issues please mention it during the UOM for discussion with machine colleagues.
CTRL:
- Today 13:00 new version for the DOOCS devices will be deployed for undulator controls. It takes 2-3 minutes.
- Follow up of PP-ODL devices: Wajid E. will right a ticket to DOOCS colleagues.
DA:
- Nothing to report.
EEE:
- Yesterday a SA2 SEPS interlock was updated and tested with XRO group and it went fine.
VAC:
- V0 closed after the SA2 loop update which was promptly re-opened.
XPD:
- Nothing to report.
DET:
- Not present.
ITDM:
- Upgrade of GPFS was finished yesterday successfully.
XRO:
- SEPS interlock update in SA2 worked well.
- One issue: communication to DOOCS was unstable in the early afternoon. Late on the day it was better.
LAS:
- Nothing to report. All fine.
UNSYS:
- At 15:30 a ZZ will be done to fix 3-way valve at SA2/cell4.
- Issues with MARWIN robot in SA3 and A. Leuschner will make a ZZ for fixing it.
SPB:
- Small issues on last THU and MON related to middle layer device and macros. DOC is informed and have recognized it. Further investigation in the particular macro that may can cause the issue.
FXE:
- Preparing for next week.
MID:
- All fine at MID.
HED:
- Nothing special; preparing next experiment.
SCS:
- Nothing to report.
SQS:
- Not present.
SXP:
- Not present. |